find the roots using simple fixed point iteration. Stop until at 15th iteration or when approximate percent relative error is below 0.05%.

3x^4-8x^3-37x^2+2x+40
1
Expert's answer
2020-10-22T17:26:28-0400
"3x^4-8x^3-37x^2+2x+40="

"=(3x^4-3x^3) -(5x^3-5x^2)-(42x^2-42x)-(40x+40)="

"=(x-1)(3x^3-5x^2-42x-40)="

"=(x-1)((3x^3+6x^2)-(11x^2+22x)-(20x+40))="

"=(x+2)(x-1)(3x^2-11x-20)="

"=(x+2)(x-1)((3x^2-15x+(4x-20))="

"=(x+2)(x-1)(x-5)(3x+4)"

"x1=-2, x2=-\dfrac{4}{3}\approx-1.333, x3=1,x4=5"

We know that there is a solution for the equation "3x^4-8x^3-37x^2+2x+40=0" in "[0,2]."


"x=\big(\dfrac{3x^4-8x^3+2x+40}{37}\big)^{1/2}"

"x_{i+1}=\big(\dfrac{3x_i^4-8x_i^3+2x_i+40}{37}\big)^{1/2}"

"\begin{matrix}
 i & x_i \\
 0 & 0\\
 1 & 1.039750 \\
 2 & 0.994489\\
 3 & 1.000742\\
 4 & 0.999900\\
 5 & 1.000014\\
 6 & 0.999998\\
 7 & 1.000000 \\
 8 & 1.000000\\
 
\end{matrix}"


"\varepsilon =\big|\dfrac{n_{i+1}-n_i}{n_i}\big|\cdot 100\%"

"\varepsilon =\big|\dfrac{1.039750-0}{0}\big|\cdot 100\%=undefined"

"\varepsilon =\big|\dfrac{0.994489-1.039750}{1.039750}\big|\cdot 100\%=4.5512\%"

"\varepsilon =\big|\dfrac{1.000742-0.994489}{0.994489}\big|\cdot 100\%=0.6248\%"

"\varepsilon =\big|\dfrac{0.999900-1.000742}{1.000742}\big|\cdot 100\%=0.0841\%"

"\varepsilon =\big|\dfrac{1.000014-0.999900}{0.999900}\big|\cdot 100\%=0.0114\%"

"\varepsilon =\big|\dfrac{0.999998-1.000014}{1.000014}\big|\cdot 100\%=0.0016\%"

"\varepsilon =\big|\dfrac{1.000000-0.999998}{0.999998}\big|\cdot 100\%=0.0002\%"

"x=1.000014"


The root is "1.000014"